OpenStudy (anonymous):

If m - n = 32 and 3m - 8n - 6 = 0, then 4m - 9n =

OpenStudy (texaschic101):

m - n = 32 m - 32 = n so plug in m - 32 in for n in this equation : 3m - 8n - 6 = 0 3m - 8(m - 32) - 6 = 0 -- distribute 3m - 8m + 256 - 6 = 0 -5m + 250 = 0 -5m = -250 m = -250/-5 m = 50 m - n = 32 50 - n = 32 50 - 32 = n 18 = n check.. 3m - 8n - 6 = 0 3(50) - 8(18) - 6 = 0 150 - 144 - 6 = 0 150 - 150 = 0 0 = 0 (correct) so we know that m = 50 and n = 18 now we sub 4m - 9n = 4(50) - 9(18) = 200 - 162 = 38